# Identifying Fairness Issues in Automatically Generated Testing Content

### v1

@@ 5 Test Results
Figure 3: F1 scores on two test sets for each proposed method. Note that for generic (short), the results were 0 on the unknown test set. Full results in Appendix E.
The previous experiments describe our attempts to identify the best-performing model for fairness classification on our validation set. Our goal is to develop a system that generalizes: for this, we evaluate the best performing of the above model types on two held-out test sets:
@@
    Out-of-domain: All samples (85) from the three held-out types: Conversations, Respond to a Written Request, and Announcements.
@@
Traditional transformer-based classification performs remarkably well, especially in generalizing to the out-of-domain data. On the in-domain data, the best performing model bert-large-cased performs on par with the best base prompting model (.571 compared to .6 F1 score). On the out-of-domain data, roberta-base performs nearly as well as the best-performing overall model, just .04 behind the generic (short) prompt with self-correction and few-shot learning. If the goal is to quickly and cheaply build a system that is applicable to a wide variety of domains, there appears to be significant value in relying on these relatively small transformer-based classification models. The Topic (data) approach is also competitive on out-of-domain data, and doesn’t even require model training: it lags only slightly behind the roberta-base model.
@@
We found significant success in our proposed self-correction mechanism. While it typically doesn’t outperform few-shot learning in isolation, the methods are naturally complementary, and the combination often yields the best-performing model.
In our experiments, we noted some issues. First, when run using too many samples or batches, the prompts tend to degrade: once the LLM makes an error and returns a prompt that doesn’t match the specifications, the run needs to be aborted. Even when the LLM sticks to the instructions, after many iterations the prompts become unwieldy and self-contradictory, and performance rapidly declines. We suggest using somewhere between 6 and 20 total samples for prompt self-correction: it’s best to not continue to make corrections indefinitely.
@@
We here report F1 score as a balance between precision and recall (for full scores, see Appendix E). Depending on the end use-case, other metrics may be more appropriate. In our case we advocate always including humans in the evaluation process to ensure that only fair content is accepted: we then value both precision (as we don’t want to unnecessarily flag excessive content as having an issue, which could reduce diversity) and recall (as we don’t want to let fairness issues through). Optimizing for recall seems reasonable, as it is likely more important to prevent fairness issues from being released, but it is critical to note that no system is perfect: even optimizing for recall, these fairness issues are likely to persist, and the models should not be used as failproof safeguards.
@@
Fine-tuned bert-base-cased 0.0 .107
bert-large-cased .116 .107
roberta-base .116 .320
Topic-based Data .289 .571
Guideline-based .182 .052
Base Prompting generic (short) 0.0 0.0
guideline (short) .226 .077
data-driven .381 .500
Self-correction generic (short) .161 .235
guideline (short) .277 .103
data-driven .271 .210
Few-shot generic (short) .212 .419
guideline (short) .445 .313
data-driven .322 .552
Few-shot + Self-correction generic (short) .387 .393
guideline (short) .445 .287
data-driven .264 .470
@@
We evaluate performance on the test set for the two subcategories annotated: Knowledge, Skill, and Ability (KSA) and Emotion (Table 3). Models built on data (the data-driven prompts (.47) and Topics from Data (.571)) perform best on the Emotion, likely due to the inclusion of negative emotional issues in the text. The data-based Topic classification model performs best in this category, highlighting the benefits of quick, untrained models. The guideline (short) prompt combined with few-shot and self-correction performs best in KSA identification (.445), highlighting the difficulty of the task.
@@ 6 Conclusions
This work delivers four key contributions: an exploration of a novel fairness detection task, a dataset of 620 samples annotated for fairness issues, evaluation of a variety of classification models for this task, including fine-tuning, topic-based approaches, and prompting, and a novel prompting strategy, which, combined with few-shot learning, achieves state-of-the-art performance on the task.
This work is aimed to explore the space of fairness and bias issues in generated content, especially in the education context. We aim to highlight the difficulties of accounting for fairness particularly in specific contexts unlikely to be accounted for by traditional model guardrails. As language model usage becomes more prevalent, the need for proper bias and fairness strategies from those training, deploying, and using these models is paramount.
@@ 7 Limitations
Our work is limited largely by the type of content evaluation and the models used. We focus specifically on a small number of item and task types that fall under very specific fairness constraints: the evaluation of the methods used specifically applies to these items under these constraints. This is apparent in the evaluation on the "unseen" item types in Section 5. Applying these methods to new item and task types, even those annotated under the same fairness guidelines, yields significantly reduced results: this is evidence that the methods and models designed work only for the specific contexts in which they are trained and developed.
Similarly, we explore a relatively small space of models and approaches. We use relatively basic prompt strategies: there exist many other approaches and improvements that are likely to be valuable that we do not evaluate. The same is true of fine-tuned models, and topic classification: we present relatively basic, well-known strategies to better understand the difficulty of our data, with the understanding that there are substantial improvements that could be applied.
8 Ethics
Content generation comes with inherent ethical concerns relating to fairness, bias, factuality, and sensitivity. Our work aims to mitigate some of these issues with regard to fairness, but it is important to consider potential issues that might arise from using large language models and other NLP technology in generating assessment content. Models may introduce subtle biases against disadvantaged groups, or produce content that appears to be factual, but is not: these are critical failures that need to be accounted for.
In practice, the generation of assessment content requires human intervention: large language model generations are not at the point where they are immune to these negative impacts, and thus for any content that goes into production, a human with relevant expertise needs to evaluate it. The methods we propose support this human intervention: these methods can remove obviously offensive content before the human even reviews, or assist in human reviews by flagging potentially harmful content.
While our dataset is unlikely to contain any content that is triggering (our framework of fairness is focused on more nuanced contexts), it must be noted that there is potential for it to be used maliciously, for example, by someone designing a system to adapt to and deceive a fairness detection systems. In releasing this data we hope to bring awareness to this issue, and better understand the potential negative impacts. Primarily, we stress that any fairness detection system should not be used in isolation or without supervision as a catchall for potential issues.
@@ Appendix A Item Types
*Announcements Ladies and gentlemen, this is John Smith from Campus Administration. We would like to remind you of our noise policies. Please refrain from loud activities in study areas and respect quiet hours in residential buildings. Violators may face disciplinary action. Thank you for your cooperation.
Table 4: Item types and annotations for fairness issues. Each has a binary annotation for containing a fairness issue and then can be also tagged as containing a KSA issue or an Emotion issue. Multiple subcategories are allowed. *Items are held out for testing, and not used for any training/evaluation.
@@ Appendix B Prompts
Ensure it is fair to for a test taker. It should not contain biased, toxic, or unethical content. It should not reflect any special skills, abilities, or culture, and it should not contain any negative emotion.
@@ Appendix C Self-Correction Prompts
After final batch of corrections
@@ Appendix D Validation Results
D Base Prompting Prompt Prec Rec F1 Generic (short) 0.0 0.0 0.0 Generic (long) 0.0 0.0 0.0 Guidelines (short) .667 .250 .364 Guidelines (long) 1.0 .042 .080 Data-driven .875 .583 .700
@@
Generic (short) 3 .917 .458 .611
5 .917 .458 .611
Guidelines (short) 3 .875 .583 .700
5 .875 .583 .700
Data-driven 3 .810 .708 .771
5 .889 .667 .762
@@
Generic+correction .500 .292 .368
Guidelines (short)+correction .900 .375 .529
Data-driven+correction .850 .708 .773
@@
Generic+correction 3 .813 .542 .650
Generic+correction 5 .789 .625 .698
Guideline+correction 3 .875 .583 .700
Guideline+correction 5 .875 .583 .700
Data-driven+correction 3 .900 .750 .818
Data-driven+correction 5 .900 .750 .818
@@ Appendix E Test Results
Fine-tuning bert-base-cased 1.0 .292 .452 .333 .056 .095
bert-large-cased .909 .417 .571 .600 .167 .261
roberta-large .900 .375 .529 .833 .278 .417
Topic-based Topics from Data .636 .292 .400 .545 .333 .414
Topics from Guidelines .714 .208 .323 .111 .056 .074
Base prompting Generic (short) 1.0 .042 .080 0.0 0.0 0.0
Guidelines (short) .625 .208 .313 .750 .167 .273
Data-driven .722 .542 .619 .462 .333 .387
Few-shot $n=3$ Generic (short) .722 .542 .619 .444 .222 .296
Guidelines (short) .929 .542 .684 .545 .333 .414
Data-driven .800 .667 .727 .500 .278 .357
Self-correction Generic (short)+correction .500 .333 .400 .286 .111 .160
Guideline (short)+correction .636 .292 .400 .333 .167 .222
Data-driven+correction .818 .375 .514 .429 .167 .240
Self-correction + few-shot $n=3$ Generic+correction .750 .500 .600 .750 .333 .462
Guideline (short)+correction .857 .500 .632 .600 .333 .429
Data-driven+correction .850 .708 .773 .500 .167 .250

### v2

@@ 5 Test Results
Figure 4: F1 scores on two test sets for each proposed method. Note that for bert-large-cased and generic (short), the scores were 0.00 on the unknown test set. Full results in Appendix E.
The previous experiments describe our attempts to identify the best-performing model for fairness classification on our validation set. Our goal is to develop a system that generalizes. For this, we evaluate the best-performing of the above model types on two held-out test sets:
@@
    Out-of-domain: All samples (66) from the two held-out types: Conversations, Respond to a Written Request.
@@
Traditional transformer-based classification performs remarkably well, especially in generalizing to the out-of-domain data. On the in-domain data, the best performing model deberta-base performs on par with the best base prompting model (0.58 compared to 0.60 F1 score), although this is a significant drop from the validation performance of 0.77, and performs quite poorly on out-of-domain data (0.20), indicating the model may overfit during training. On the out-of-domain data, roberta-base performs nearly as well as the best-performing overall model, just 0.04 behind the generic (short) prompt with self-correction and few-shot learning. If the goal is to quickly and cheaply build a system that is applicable to a wide variety of domains, there appears to be significant value in relying on these relatively small transformer-based classification models. The Topic (data) approach is also competitive on out-of-domain data, and does not even require model training; it lags only slightly behind the roberta-base model.
@@
We found significant success in our proposed self-correction mechanism. While it typically does not outperform few-shot learning in isolation, the methods are naturally complementary, and the combination often yields the best-performing model. In examining the models’ self-corrections, we find that when asked to become more restrictive, the model tends to add sentences with new constraints, which nicely reflect the issue that was missed. When asked to become less restrictive, the model tends to add hedges to currently existing constraints.
In our experiments, we noted some issues. First, when run using too many samples or batches, the prompts tend to degrade: once the LLM makes an error and returns a prompt that doesn’t match the specifications, the run needs to be aborted. Even when the LLM sticks to the instructions, after many iterations the prompts become unwieldy and self-contradictory, and performance rapidly declines. We suggest using somewhere between six and 20 total samples for prompt self-correction; it is best to avoid making corrections indefinitely.
@@
We here report F1 score as a balance between precision and recall. (For full scores, see Appendix E.) Depending on the end use case, other metrics may be more appropriate. In our case, we advocate for always including humans in the evaluation process to ensure that only fair content is accepted. We then value both precision (as we do not want to excessively flag content for fairness issues, which could reduce diversity) and recall (as we do not want to let fairness issues through). Optimizing for recall seems reasonable, as it is likely more important to prevent fairness issues from being released, but it is critical to note that no system is perfect: even optimizing for recall, these fairness issues are likely to persist, and the models should not be used as failproof safeguards.
@@
Fine-tuned bert-base-cased 0.07 0.57
bert-large-cased 0.00 0.00
roberta-base 0.06 0.56
deberta-base 0.08 0.75
Topic-based Data 0.26 0.59
Guideline-based 0.20 0.06
Base Prompting generic (short) 0.00 0.00
guideline (short) 0.29 0.09
data-driven 0.47 0.50
Self-correction generic (short) 0.35 0.30
guideline (short) 0.35 0.27
data-driven 0.47 0.41
Few-shot generic (short) 0.18 0.24
guideline (short) 0.30 0.24
data-driven 0.36 0.56
Few-shot + Self-correction generic (short) 0.18 0.21
guideline (short) 0.23 0.21
data-driven 0.24 0.59
@@
We evaluate performance on the test set for the two subcategories: Knowledge, Skill, and Ability (KSA) and Emotion (Table 3). The deberta-base model performs exceptionally well on the KSA subcategory, capturing 75% of the fairness-flagged samples. Data-based methods (the data-driven prompts (0.59) and Topics from Data (0.59)) also perform well, likely due to the inclusion of negative emotional issues in the text. They perform much worse on KSA classification, although the data-driven prompts still yield the best performance (0.47): KSA-related issues are especially difficult as they generally involve only specific knowledge, and would not normally be considered fairness issues in other contexts.
@@ 6 Conclusions
This work delivers four key contributions: an exploration of a novel fairness detection task, a dataset of 601 samples annotated for fairness issues, evaluation of a variety of classification models for this task, including fine-tuning, topic-based approaches, and prompting, and a novel prompting strategy, which, combined with few-shot learning, achieves state-of-the-art performance on the task.
This work is aimed to explore the space of fairness and bias issues in generated content, especially in the education context. We aim to highlight the difficulties of accounting for fairness, particularly in specific contexts unlikely to be accounted for by traditional model guardrails. As language model usage becomes more prevalent, the need for proper bias and fairness strategies from people training, deploying, and using these models is paramount.
7 Ethics
Content generation comes with inherent ethical concerns relating to fairness, bias, factuality, and sensitivity. Our work aims to mitigate these issues with regard to fairness, but it is important to consider potential issues that might arise from using LLMs and other NLP technology in generating assessment content. Models may introduce subtle biases against disadvantaged groups, or produce content that appears to be factual, but is not. These are critical failures that need to be accounted for.
In practice, the generation of assessment content requires human intervention: large language model generations are not at the point where they are immune to these negative impacts, and thus for any content that goes into production, a human with relevant expertise needs to evaluate it. The methods we propose support this human intervention, as they can remove obviously offensive content before the human review stage, or assist in human reviews by flagging potentially harmful content.
While our dataset is unlikely to contain any content that is triggering (our framework of fairness is focused on more nuanced contexts), it must be noted that there is potential for it to be used maliciously; for example, by someone designing a system to adapt to and deceive a fairness detection system. In releasing this data, we hope to bring awareness to this issue and better understand the potential negative impacts. Primarily, we stress that any fairness detection system should not be used in isolation or without supervision as a catchall for potential issues.
@@ 8 Limitations
Our work is limited largely by the type of content evaluation and the models used. We focus on a small number of item and task types that fall under very specific fairness constraints: the evaluation of the methods used specifically applies to these items under these constraints. This is apparent in the evaluation on the "unseen" item types in Section 5. Applying these methods to new item and task types, even those annotated under the same fairness guidelines, yields significantly reduced results. This is evidence that the methods and models we designed work only for the specific contexts in which they are trained and developed.
Similarly, we explore a small space of models and approaches. We use relatively basic prompt strategies; there exist many other approaches and improvements that are likely to be valuable that we do not evaluate. The same is true of fine-tuned models and topic classification. We present relatively basic, well-known strategies to better understand the difficulty of our data, with the understanding that there are substantial improvements that could be applied.
@@ Appendix A Item Types
Table 4: Item types and annotations for fairness issues. Each has a binary annotation for containing a fairness issue and then can be also tagged as containing a KSA issue or an Emotion issue. Multiple subcategories are allowed. *Items are held out for testing, and not used for any training/validation.
@@ Appendix B Prompts
Ensure it is fair to a test taker. It should not contain biased, toxic, or unethical content. It should not reflect any special skills, abilities, or culture, and it should not contain any negative emotion.
@@ Appendix C Self-Correction Prompts
After a final batch of corrections
@@ Appendix D Validation Results
D Base Prompting Prompt Prec Rec F1 Generic (short) 0.00 0.00 0.00 Generic (long) 0.00 0.00 0.00 Guidelines (short) 0.67 0.25 0.36 Guidelines (long) 1.00 0.04 0.08 Data-driven 0.88 0.58 0.70
@@
Generic (short) 3 0.92 0.46 0.61
5 0.92 0.46 0.61
Guidelines (short) 3 0.88 0.58 0.70
5 0.88 0.58 0.70
Data-driven 3 0.81 0.71 0.77
5 0.89 0.67 0.76
@@
Generic+correction 0.50 0.29 0.37
Guidelines (short)+correction 0.90 0.38 0.53
Data-driven+correction 0.85 0.71 0.77
@@
Generic+correction 3 0.81 0.54 0.65
Generic+correction 5 0.79 0.63 0.70
Guideline+correction 3 0.88 0.58 0.70
Guideline+correction 5 0.88 0.58 0.70
Data-driven+correction 3 0.90 0.75 0.82
Data-driven+correction 5 0.90 0.75 0.82
@@ Appendix E Test Results
Fine-tuning bert-base-cased 1.00 0.29 0.45 0.75 0.20 0.32
bert-large-cased 0.91 0.42 0.57 0.00 0.00 0.00
roberta-large 0.90 0.38 0.53 0.67 0.13 0.22
deberta-base 1.00 0.42 0.58 1.00 0.13 0.24
Topic-based Topics from Data 0.64 0.29 0.40 0.57 0.27 0.36
Topics from Guidelines 0.71 0.21 0.32 0.14 0.07 0.09
Base prompting Generic (short) 1.00 0.04 0.08 0.00 0.00 0.00
Guidelines (short) 0.63 0.21 0.31 0.50 0.27 0.35
Data-driven 0.72 0.54 0.62 0.47 0.47 0.47
Few-shot $n=3$ Generic (short) 0.72 0.54 0.62 1.00 0.67 0.13
Guidelines (short) 0.93 0.54 0.68 0.50 0.13 0.21
Data-driven 0.80 0.67 0.73 1.00 0.33 0.50
Self-correction Generic (short)+correction 0.50 0.33 0.40 0.56 0.33 0.42
Guideline (short)+correction 0.64 0.29 0.40 0.50 0.47 0.48
Data-driven+correction 0.82 0.38 0.51 0.56 0.33 0.42
Self-correction + few-shot $n=3$ Generic+correction 0.75 0.50 0.60 1.00 0.07 0.13
Guideline (short)+correction 0.86 0.50 0.63 0.33 0.07 0.11
Data-driven+correction 0.85 0.71 0.77 0.67 0.13 0.22